OpenStudy (anonymous):

Compute the following Bernoulli Trials: 1.) 3 successes in 5 trials with p = 2/7? 2.) 0 successes in 4 trials with p = 1/6?

OpenStudy (anonymous):

do you know the formula for this?

OpenStudy (anonymous):

n = 5, k = 3, p = 2/7, 1-p= 5/7

OpenStudy (anonymous):

yes but I cant get it to work for some reason I have tried numerous times and cant get the correct solution

OpenStudy (anonymous):

\[P(k=3)=\dbinom{5}{3}(\frac{2}{7})^3\times( \frac{5}{7})^2\]

OpenStudy (anonymous):

three sucesses explains the \[(\frac{2}{7})^3\] 2 failure explain \[(\frac{5}{7})^2\] and there are \[\dbinom{5}{3}=10\] ways to arrange the successes and failures

OpenStudy (anonymous):

here is what i got http://www.wolframalpha.com/input/?i=10%282%2F7%29^3*%285%2F7%29^2

OpenStudy (anonymous):

second one is easier, all failures is \[(\frac{5}{6})^4\]

OpenStudy (anonymous):

you got this ok?

OpenStudy (anonymous):

I think so it is a little clearer thank you so much for your help

OpenStudy (anonymous):

yw